5. In a class test, the sum of Shefali’s marks in Mathematics and English is 30. Had she got 2 marks more in Mathematics and 3 marks less in English, the product of their marks would have been 210. Find her marks in the two subjects.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let us say, the marks of Shefali in Maths be x. Then, the marks in English will be 30 – x. As per the given question, (x + 2)(30 – x – 3) = 210 (x + 2)(27 – x) = 210 ⇒ -x2 + 25x + 54 = 210 ⇒ x2 – 25x + 156 = 0 ⇒ x2 – 12x – 13x + 156 = 0 ⇒ x(x – 12) -13(x – 12) = 0 ⇒ (x – 12)(x – 13) = 0 ⇒ x = 12, 13 Therefore, if the marks in Maths are 12, then marks in English will be 30 – 12 = 18 and the marks in Maths are 13, then marks in English will be 30 – 13 = 17.